Circular No. 8274

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2003 SQ_{215} (NEAT-LONEOS) An apparently asteroidal object reported independently by the NEAT

and LONEOS projects (discovery observations below, from MPS 86352) has been found to show a nonstellar appearance in individual 30-s R-band images taken by A. Fitzsimmons and C. Snodgrass, Queen’s University of Belfast, and O. Hainaut, European Southern Observatory (ESO), on 2004 Jan. 19.0 UT at the ESO 3.6-m New Technology Telescope (+ SUSI-2 camera). Fitzsimmons adds that co-addition of the frames shows an asymmetric coma of total mag 20.3 extending 1′′.7 in p.a. 130^{o}.

2003 UT α_{2000} δ_{2000} Mag. Observer
Sept.24.18377 21^{h}42^{m}42.^{s}67 −5^{o}13′24′′.2 18.9 NEAT
27.16490 21 41 46.21 −5 22 30.6 18.7 LONEOS

Additional astrometry (see also MPEC 2004-A38), the following orbital elements, and an ephemeris appear on MPEC 2004-B35.

Epoch = 2004 Mar. 16.0 TT

T = 2004 Mar. 24.2293 TT  ω = 137.^{o}3029
e = 0.581259  Ω = 257.6406 }2000.0
q = 2.304226 AU  i = 5.5459

a = 5.502741 AU n^{o} = 0.0763546 P = 12.908 years

SUPERNOVA 2004L IN MCG +03-27-38 Further to IAUC 8273, H. Pugh and W. Li report the LOSS discovery,

on KAIT images taken on Jan. 21.4 (mag ∼ 16.9) and 22.4 UT (mag ∼ 16.8), of an apparent supernova located at α = 10^{h}27^{m}04^{s}.11, δ = +16^{o}01′07′′.3 (J2000), which is 2′′.5 east and 3′′.1 north of the nucleus of MCG +03-27-38. A KAIT image on 2003 Nov. 25.4 showed nothing at this position (limiting mag ∼ 19.0).

DO DRACONIS A. Price, AAVSO, reports that DO Dra (cf. IAUC 7260) is in outburst

for the first time since 2000 Nov. 15 (when it peaked at m_{v} = 10.1 and faded to quiescence in 4 days), as indicated by the following visual magnitude estimates: Jan. 23.229 UT, 14.5 (M. Simonsen, Imlay City, MI); 23.733, 11.4 (W. Kriebel, Osterwaal, Germany); 23.758, 11.0 (C. Otten, Kinrooi, Belgium); 23.808, 11.5 (W. Renz, Karlsruhe, Germany).

2004 January 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green